OpenStudy (anonymous):

Amy has a triangular flag. She wants to cut out a similar flag that has dimensions one-sixth of the original flag. If the area of the original flag is 2160 square inches, what is the area of the new flag that Amy cuts out?

Directrix (directrix):

The two flags will be similar. If two polygons are similar, the ratio of their areas is equal to the square of the scale factor of the two polygons. (1/6)^2 = A/2160 where A is the area of the smaller flag 1/36 = A/2160 36A = 2160 A = 60 square inches